# Chaos and thermalization in Clifford–Floquet dynamics

### v1

@@ 6. Discussion
In this note we generalized the results of [CQCA2] to Clifford QCAs in general dimensions and an arbitrary number of qubits per site. We showed that soliton-free Clifford QCAs (i.e. those Clifford QCAs which do not preserve any local quantity up to translation) thermalize a large class of states, including all SRE states sufficiently close to the equilibrium state and states obtained from these by local von Neumann measurements. Moreover, our numerical simulations indicate that soliton-free Clifford QCAs thermalize a larger class of SRE states which are quite far from equilibrium and are not covered by our rigorous results. All of this evidence supports the identification of soliton-free Clifford QCAs as paradigmatic examples of chaotic deterministic quantum dynamics.
@@
Finally, although we stated our results for systems of qubits, identical results hold for systems of qudits.

### v2

@@ 6. Discussion
In this paper, we examined quantum chaotic behavior arising in Floquet quantum spin systems without disorder. This is to be contrasted with quantum spin systems driven by random unitary circuits (see [6] for a review) or spatially random circuits [17, 5]. We generalized the results of [7] to Clifford QCAs in general dimensions and an arbitrary number of qubits per site. For simplicity, we stated our results for systems of qubits, but identical results hold for systems of qudits. We showed that soliton-free Clifford QCAs (i.e. those Clifford QCAs which do not preserve any local quantity up to translation) thermalize a large class of states, including all SRE states sufficiently close to the equilibrium state and states obtained from these by local von Neumann measurements. Moreover, our numerical simulations indicate that soliton-free Clifford QCAs thermalize a larger class of SRE states which are quite far from equilibrium and are not covered by our rigorous results. All of this evidence supports the identification of soliton-free Clifford QCAs as examples of chaotic deterministic quantum dynamics.